A particle moves in a circle of radius 20 cm with a linear speed of 10 m/s find the angular velocity

Grade:11

2 Answers

Arun
6 years ago
\omega =\frac{v}{r}=\frac{10}{20}=0.5 rad/secDear student,The angular velocity is given by:as we knowv = r* wWhere r = radius, v = linear velocity, w = angular velocityHope it helps
 
anjace
6 years ago
w=v/r
10/20=o.s rad/s............................................................................................................................................................................................................................